#117181
Express the roots of the equation z 3 − α 3 = 0 in terms of α and w, where w is a complex cube root of unity. Use your answer to find the roots of the following equations in the form a + ib. (a) z 3 − 27 = 0 (b) z 3 + 8 = 0 (c) z 3 − i = 0.
1
Expert's answer
2020-05-20T16:42:57-0400

For the equation "z^{3}=a^{3}" , we have that

"z^{3}=a^{3} * 1=a^{3}*e^{(2\pi)i}=a^{3}(cos 2\pi + isin 2\pi)"

Since "sin 2\pi , cos 2\pi" are periodic functions of the period "2\pi k , k=0,1,2" , then

"z_{k}^{3}=a^{3}(cos( 2\pi+ 2\pi k)+ isin ( 2\pi+ 2\pi k))"

thus "z_{k}=a(cos( 2\pi+ 2\pi k)+ isin ( 2\pi+ 2\pi k))^{1/3}"

using De Moivre's formula, we have

"z_{k}=a(cos( 2\pi+ 2\pi k)/3+ isin ( 2\pi+ 2\pi k)/3)" , where "k=0,1,2"

when k=0 , we have that the first root "z_{0}=aw , w=-1/2+(\sqrt3/2)i"

when k=1 , we have that the second root "z_{1}=aw^{2}=a(-1/2-(\sqrt3/2)i)"

when k=2 , we have that the third root "z_{2}=a"

### To solve the problem (a) we put a=3 , then the roots are "-3/2+(3\sqrt3/2)i" , "-3/2-(3\sqrt3/2)i" , 3

### To solve the problem (b) , we put "a=-2" , thus the roots are "1-\sqrt3i , 1+\sqrt3i , -2"

### But the problem (c) is different, so we solve it with different steps

"z^{3}=i=cos \pi/2+ isin \pi/2" , then "z^{3} = cos( \pi/2+ 2\pi k)+ isin ( \pi/2+ 2\pi k)"

where ", k=0,1,2"

thus the roots are "z_{k} = cos(( \pi/2+ 2\pi k)/3)+ isin ( ( \pi/2+ 2\pi k)/3 )"

when k=0 , "z_{0} = cos( \pi/6)+ isin ( \pi/6)"

when k=1 , "z_{1} = cos( 5\pi/6)+ isin ( 5\pi/6 )"

when k=2 , "z_{2} = cos(9\pi/6)+ i sin ( 9\pi/6 )"